A piston-cylinder device contains 0.5 m3 air at 100 kPa and 1000 K. Because the outside is cooler (at 300 K), the hot air starts losing heat to the surroundings and the volume of air starts decreasing. Over a certain period of time the volume is observed to shrink by 50%.

The final temperature is lower or higher?

Determine the boundary work transfer (absolute value). in KJ

was W positive and Q negative? or what?
Answer #1


the final temparature is lower than the 1000 k

and higher than the 300 k

boundary work transfer is W = P*(V2-V1) = 100*1000*(0.25-0.5) = -25 kJ

absolute value is 25 kJ

W is negative and Q is positive